OpenStudy (anonymous):

how would you find an equivalent expression to 3c(1/3d-9)-7(c+1)+d(c+4)

OpenStudy (anonymous):

what do you have so far

OpenStudy (anonymous):

we tried distributing but we havnt gotten any of the answer choices

OpenStudy (anonymous):

need to see the answer choices

OpenStudy (anonymous):

A) 2cd-34c+4d-7 B) 2cd-7c-4 C) 2cd+34c+4d+7 D) 2cd+8c+4

OpenStudy (anonymous):

could you explain how you set that up? we havnt learned much this year and we have an EOCT on tuesday.

OpenStudy (anonymous):

hold on im dumb

OpenStudy (anonymous):

haha ok thank you

OpenStudy (anonymous):

is it 1/3d-9 or 1/(3d-9)

OpenStudy (anonymous):

3c(1/3d-9)

OpenStudy (anonymous):

\[3 \frac{ 1 }{ 3d-9 } - 7(c+1)+d(c+4)\]

OpenStudy (anonymous):

\[=\frac{ 1 }{ d-3 } -7c -7 +cd + 4d\]

OpenStudy (anonymous):

\[= \frac{1+ (d-3)(-7c-7+cd+4d) }{ (d-3) }\]

OpenStudy (anonymous):

oook thanks

OpenStudy (anonymous):

id keep going but the problem looks weird for some reason

OpenStudy (anonymous):

yeah thanks tho